Determine the values of the random variables in each of the following distributions.




1. Two coins are tossed. Let H be the number of tails that occur. Determine the values of the random variable H.




of the random variable K representing the number of Koreans.

1
Expert's answer
2022-01-19T18:08:24-0500

1) "P(H=0)=\frac{1}{2}*\frac{1}{2}=\frac{1}{4}."

"P(H=1)=\frac{1}{2}*\frac{1}{2}*2=\frac{1}{2}."

"P(H=2)=\frac{1}{2}*\frac{1}{2}=\frac{1}{4}."


2) "P(K=0)=0."

"P(K=1)=\frac{C_2^2*C_4^1}{C_6^3}=\frac{1}{5}."

"P(K=2)=\frac{C_2^1*C_4^2}{C_6^3}=\frac{3}{5}."

"P(K=3)=\frac{C_2^0*C_4^3}{C_6^3}=\frac{1}{5}."